# GalaxyGenius: A Mock Galaxy Image Generator for Various Telescopes from Hydrodynamical Simulations

### v1

@@ 2.2 Data preprocessing
Several critical parameters influence the output of SKIRT. These parameters include simulation mode, dust model, wavelength range, number of photon packets and some instrument-specific configurations such as field of view, number of spatial pixels, observing directions, and distance. Since interstellar dust primarily impacts the infrared wavelength range, we include dust in the radiative transfer procedures only for observations that extend into the infrared. In such cases, we adopt the DustEmission mode to account for dust absorption, scattering, and thermal re-emission. For modeling dust grains, SKIRT offers several options, including dust model introduced in Zubko et al. (2004) (ZubkoDustMix), Draine & Li (2007) (DraineLiDustMix) and Jones et al. (2017) (ThemisDustMix). We adopt a fixed dust grain model, ZubkoDustMix, in every location of the galaxy. This model comprises three dust components: silicate, graphite, and polycyclic aromatic hydrocarbons (PAH), with size distributions defined in 15 bins. We employ hierarchical octree with refinement level from 6 to 10 to re-grid the dust density to increase the speed of radiative transfer for dust, instead of using the native Voronoi grid in TNG simulation (Baes et al., 2024). Further details on other dust models and their effects on generating galaxies are provided in Section 4.1. While for observations that do not reach infrared, we utilize NoMedium mode, which is computationally more efficient as it exclude dust elements and focuses solely on emission from primary sources. The wavelength range should be set to encompass the wavelength of all the considered filters, taking into account the effect of redshift. The number of photon packets in Monte Carlo simulation plays a crucial role in determining the signal-to-noise ratio (SNR) of the output and the computational cost. To strike a balance between SNR and efficiency, we use a fixed value of photon packets for all simulations. This choice ensures high-quality results while maintaining reasonable computational demands. Regarding instrument-specific configurations, the field of view can be freely adjusted, depending upon the aperture size of considered instrument. Consequently, the outputs including data cubes and SEDs will be limited to this specific size. The number of spatial pixels should be set to correspond to the pixel scale of the specific filter. Otherwise, rescaling is necessary to generate mock images. It is advisable to set observing directions with caution, as the memory usage significantly increases with the number of observing directions. Observing distance can be calculated based on redshift if a flat universe is assumed to account for the cosmological surface brightness dimming. In contrast, for a local universe, the distance should be explicitly specified.
@@
The snapshot and subhalo data of IllustrisTNG, from which particles are retrieved, are of an exceptionally large size. Fortunately, a web-based interface (API) 333https://www.tng-project.org/data/docs/api/ can be utilized to search, extract, visualize, and analyze data without the necessity of acquiring the entire simulation data. The utilization of this web-API is also implemented in GalaxyGenius to facilitate galaxy generation.
@@ 2.3 Data cube generation
The radiative transfer procedure is executed using the particle files and the SKIRT execution file generated during the preprocessing stage mentioned above. This process assigns each particle a SED based on its properties and simulates the propagation of light through the ISM of the galaxy, accounting for absorption, scattering, and emission. The memory consumption during this step is primarily influenced by two factors: spatial and wavelength resolution and the number of viewing angles. Higher resolution and more viewing angles increase the computational and memory requirements significantly.
@@ 2.4 Mock observation
After acquiring the data cubes, we can calculate the bandpass images, given the throughput of each filter for a particular instrument. All calculations are performed in units of electron counts, enabling straightforward incorporation of PSF effects and instrumental and background noises. The electron counts for each pixel can be determined by providing the exposure time , the number of exposures , the effective area of the telescope , and the pixel scale of the CCD camera as follows (see Section 9.2 in Ryon (2023)):
where and are Planck constant and the speed of light respectively. is the throughput of filter and is flux measured per wavelength, converted from flux measured per frequency in unit of employed by SKIRT using following equation:
The PSF describes how a point source of light, such as a star, is distributed across a detector, and is a fundamental characteristic of imaging system, as it determines the system’s ability to resolve fine details. The PSF size is influenced by fundamental physical limits as well as practical effects like atmospheric turbulence, optical aberrations, and detector effects. In ground-based observations, the most dominant factor is atmospheric turbulence, while in space-based ones, detector effects play a larger role. Furthermore, the PSF exhibits variations across the position of the CCD and observational pointing. Consequently, for simplicity, in our galaxy generation for specific instrument, we only consider the PSF kernels at an arbitrary position and convolve the ideal images with these kernels. Note that when convolving with the model PSF, the pixel sizes of the image and the kernel should be the same.
@@
where indicates the pixel of image on which the Poisson components are added. And is the mean of expected electron counts in pixel, calculated as:
where is the pixels of image with PSF effects already applied. And is the noise term that includes the contribution of two components, sky background noise and dark current noise:
where and are sky background and dark current both in unit of electron count rate (). Between them, dark current comes from thermal activity within CCD, even in the absence of light, while sky background can be calculated by (see Section 9.4 in Ryon (2023)):
where is the brightness of the sky background in Å-1. The sky brightness is contributed by several components, such as airglow, zodiacal light, light pollution, terrestrial background (Ryon, 2023). These components can vary depending on factors like calendar day, observational pointing, orbit or site position, and atmospheric conditions. Therefore, we only consider brightness curves under arbitrary conditions for specific observational campaign.
Subsequently, readout noise is introduced, which results from the random fluctuations during the process of reading out the charge from a CCD. This noise is modeled as Gaussian distribution with a mean of zero and a standard deviation corresponding to the readout level , and is correlated with the number of exposures :
where represents the pixel of image with the readout noise added. Next, we subtract the mean level of sky background and dark current calculated in Equation 9:
Finally, we obtain the galaxy images expected to be observed by specific instrument. The image unit in electron counts can be converted back to flux per frequency in . This procedure of including both the instrumental and sky background contributions ensures that the simulated data accurately represents the noisy conditions encountered in real-world astronomical observations. Alternatively, the overall noise level for the galaxy images can also be determined from the limiting magnitudes for a specific band in a given survey (Fortuni et al., 2023; Bottrell et al., 2019, 2024; Martin et al., 2022; Wilkinson et al., 2024; Euclid Collaboration et al., 2023). This approach is also implemented and detailed in Appendix C. Depending on the availability of certain survey parameters, either this approach or the one described above can be utilized.
@@ 3 Sample Outputs
In this section, we showcase the generation capabilities of GalaxyGenius for a diverse range of photometric surveys conducted both in space and on ground. These surveys encompass notable ongoing and planned missions such as CSST, HST, JWST, Euclid, Roman, and HSC. Subsequently, we demonstrate the galaxies generated from EAGLE simulations.
@@ 3.1 CSST
The Chinese Space Station Telescope (CSST, Zhan (2018); Gong et al. (2019)) is a 2-meter aperture space telescope, planned to be launched at 2027. This telescope is equiped with cutting-edge instruments, including a spectrograph and a photometric system, enabling multi-wavelength observations and spectroscopic studies. The photometric system loads seven filters, and covering from near-ultraviolet to near-infrared wavelength range. The pixel scale for each filter is uniform of .
@@
Since the filters can barely reach near-infrared band, we exclude the dust components and employ NoMedium simulation framework. The wavelength range in rest-frame is set to be from 0.1 to 1.2 micron to account for all filters. And the wavelength grid is configured as linear with 300 bins. The exposure time is considered as 150 s, with exposure time as 4 times for and bands, and 2 for other bands according to the focal plane design. The sky backgrounds are calculated as and based on the high sky background emission including earthshine and zodiacal light in optical wavelength coverage at earth orbit provided by HST 444https://hst-docs.stsci.edu/acsihb/chapter-9-exposure-time-calculations/9-7-tabular-sky-backgrounds. Other necessary instrumental parameters employed in mock observation are outlined in Appendix A. We set the viewing angle, inclination and azimuth, to be both 0, meaning that the line of sight is in parallel to the -axis. Additionally, for spiral galaxy, the viewing angles are set to be face-on and edge-on. These angles are calculated based on the total angular momentum for star particles within 30 kpc sphere (McAlpine et al., 2016).
@@
Three subhalos are presented as examples, and their properties are tabulated in Table 2, including stellar mass, star formation rate, and the number of distinct particles, as discussed in Section 2.2. The ideal SEDs generated from SKIRT in redshifted wavelength range are illustrated in Figure 2, while the galaxy images in seven bands output by postprocess procedure introduced in Section 2.4 are depicted in Figure 3. The images are presented in logarithmic scale, and to mitigate the occurence of NaN values and enhance visual clarity, they are clipped at .
Subhalo 0 is an elliptical galaxy with a low SFR and a substantial number of BC03 particles. In contrast, Subhalo 31 is an irregular galaxy characterized by a high SFR and a large number of particles belonging to MAPPINGS-III SED family, which are indicative of star-forming regions. The median panel of Figure 2 exhibits low continuum and several pronounced emission lines, such as H 6563, [SIII] and Pa 10941 from low to high wavelength. Subhalo 253881, on the other hand, is a spiral galaxy with a high SFR characterized by several of aforementioned emission lines. The continuum exhibits similar shape to Subhalo 0, primarily due to the dominance of BC03 particles.
@@ 3.2 Euclid
The Euclid mission, developed by the European Space Agency (ESA), is a 1.2-meter space-based observatory designed to explore the nature of dark energy and dark matter by mapping the geometry of the Universe (Euclid Collaboration et al., 2024d). Lauched in 2023, Euclid operates at the second Lagrange point (L2) and is optimized for both visible and near-infrared observations.
@@
Euclid’s payload includes two main instruments: VISible Imaging Channel (VIS, Euclid Collaboration et al. (2024a)) and Near Infrared Spectrometer and Photometer (NISP, Euclid Collaboration et al. (2024c)). VIS covers the optical wavelength range from 500 to 900 nanometers with a pixel scale of , providing high-resolution, wide-field imaging. This instrument is designed for precise measurement of galaxy shapes, essential for weak gravitational lensing studies. While NISP operates in the near-infrared range of 920 to 2000 nm, including a photometric mode for imaging in three broad filters (Y, J, H) with pixel scales of , and a spectroscopic mode using a slitless grism. The spectroscopic mode is employed to measure redshifts of galaxies, rendering baryon acoustic oscillation (BAO) studies.
@@
We employ similar parameters to CSST mentioned in Section 3.1 to perform the simulation, and to cover the optical and near-infrared bands, the wavelength range are set to be 0.5 to 5 micron. Furthermore, DustEmission mode is activated to account for the radiative transfer process in infrared bands. We employ dust recipe as Camps et al. (2016) (Equation 2) and dust model as Zubko et al. (2004) described in Section 2.2. The exposure time are configured as 600 s with one exposure. The PSFs are modeled using a Gaussian distribution with full width at half maximum (FHWM) as and for VIS, Y, J and H bands respectively (Euclid Collaboration et al., 2024a, c). The sky backgrounds are calculated as 0.7162, 3.6863, 4.4556 and 3.8228 for these bands based on background models 555https://irsa.ipac.caltech.edu/applications/BackgroundModel/ considering that the telescope orbits at L2 point. For other necessary instrumental parameters, please consult Appendix A.
The SEDs in the near-infrared are presented in Figure 2. The Subhalo 0 remains unaffected by the presence of dust elements because of only few MAPPING-III particles. And it is evident that when dust is included, the fluxes for Subhalo 31 experience a decrease compared to those observed in optical bands. Note that this decrease does not imply that the NoMedium simulation for optical wavelength is incorrect. Since the continuum flux is low, the decreased amplitude has a negligible impact on galaxy images or photometric measurements. On the other hand, Subhalo 253881 exhibits only diminution of emission lines at face-on direction, while at edge-on direction, the fluxes are decreased as expected since light travels denser dust regions compared to observed at face-on. These distinct behaviors conclude that the dust elements primarily suppress emissions from star-forming regions. Correspondingly, the galaxy images for the VIS and NISP bands are depicted in Figure 4. The near-infrared images reveal extended structures in comparison to the observations at optical bands.
@@ 3.3 HST WFC3
The Hubble Space Telescope (HST; Freedman et al. (2001); Momcheva et al. (2016)) is a 2.4-meter optical and ultraviolet space observatory launched in 1990, orbiting Earth at an altitude of approximately 540 km. HST operates across ultraviolet (115-320 nm), visible (320-760 nm), and near-infrared (760-2500 nm) wavelength ranges, providing high-resolution imaging and spectroscopy free from atmospheric distortion. The telescope’s pointing precision, fine guidance sensors, and advanced suite of instruments enable observations of faint and distance objects with unprecedented clarity.
The Wide Field Camera 3 (WFC3, Momcheva et al. (2016); Windhorst et al. (2011); O’Connor et al. (2024)), installed during the fifth servicing mission in 2009, is one of HST’s most versatile instruments. It features two independent optical channels: UVIS and IR channel. UVIS operates in 200-1000 nm wavelength range with a pixel CCD detector, optimized for ultraviolet and visible light imaging. This channel has a pixel scale of and a field of view of approximately . While the IR channel covers 800-1700 nm with a HgCdTe array detector, with a pixel scale of , providing a field of view of . For more detailed information on instrumental design, please refer to WFC3 official website 666https://www.stsci.edu/hst/instrumentation/wfc3/instrument-design/.
The PSF for each filter is taken from the standard PSFs described in (Anderson, 2016). The instrumental and sky background noises are calculated using the instrumental parameters and taking into account the operation of HST at Earth orbit. As an example, we consider several bands F275W, F390W and F814W from UVIS channels and F105W, F110W, F125W and F160W from IR channels to illustrate the generation of galaxies from the TNG. The exposure time is similarly assumed to be 600 s with one exposure. Galaxy images in considered bands are depicted in Figure 5.

### v2

@@ 2.2 Data preprocessing
Several critical parameters influence the output of SKIRT. These parameters include simulation mode, dust model, wavelength range, number of photon packets, and some instrument-specific configurations such as field of view, number of spatial pixels, observing directions, and distance. The commonly utilized simulation modes include NoMedium, ExtinctionOnly, and DustEmission. These modes represent ideal scenarios with no dust, concentrate solely on extinction and scattering effects, and consider extinction and scattering along with dust emission, respectively. Given that dust emission primarily affects the IR wavelength range (Draine, 2003; Meny et al., 2007; da Cunha et al., 2008; Massa et al., 2014; Decleir et al., 2022), we employ ExtinctionOnly and DustEmission modes for optical and IR ranges, respectively. The NoMedium mode is preserved for testing purposes due to its rapid execution by considering the emissions solely from primary sources. For modeling dust grains, SKIRT offers several options, including the dust model introduced in Zubko et al. (2004) (ZubkoDustMix), Draine & Li (2007) (DraineLiDustMix), and Jones et al. (2017) (ThemisDustMix). We adopted a fixed dust grain model, ZubkoDustMix, in every location of the galaxy. This model comprises three dust components: silicate, graphite, and polycyclic aromatic hydrocarbons (PAHs), with size distributions defined in 15 bins. We employed hierarchical octree with refinement level from 6 to 10 to re-grid the dust density to increase the speed of radiative transfer for dust, instead of using the native Voronoi grid in TNG simulation (Baes et al., 2024). Further details on other dust models and their effects on generating galaxies are provided in Section 4.1. The wavelength range should be set to encompass the wavelength of all the considered filters, taking into account the effect of redshift. The number of photon packets in MC simulation plays a crucial role in determining the signal-to-noise ratio (S/N) of the output and the computational cost. To strike a balance between S/N and efficiency, we used a fixed value of photon packets for all simulations. This choice ensures high-quality results while maintaining reasonable computational demands. Regarding instrument-specific configurations, the field of view can be freely adjusted, depending upon the aperture size of considered instrument. Consequently, the outputs including data cubes and SEDs will be limited to this specific size. The number of spatial pixels should be set to correspond to the pixel scale of the specific filter. Otherwise, rescaling is necessary to generate mock images. It is advisable to set observing directions with caution, as the memory usage significantly increases with the number of observing directions. Observing distance can be calculated based on redshift if a flat universe is assumed to account for the cosmological surface brightness dimming. In contrast, for a local universe, the distance should be explicitly specified.
@@
The snapshot and subhalo data of IllustrisTNG, from which the particles were retrieved, are of an exceptionally large size. Fortunately, a web-based interface (API) 333https://www.tng-project.org/data/docs/api/ can be utilized to search, extract, visualize, and analyze data without the necessity of acquiring the entire simulation data. The utilization of this web-API was also implemented in GalaxyGenius to facilitate galaxy generation.
@@ 2.3 Data cube generation
The radiative transfer procedure is executed using the particle files and the SKIRT execution file generated during the preprocessing stage mentioned above. This process assigns each particle a SED based on its properties and simulates the propagation of light through the ISM of the galaxy, accounting for absorption, scattering, and emission. The memory consumption during this step is primarily influenced by two factors: the spatial and wavelength resolution and the number of viewing angles. Higher resolution and more viewing angles increase the computational and memory requirements significantly.
@@ 2.4 Mock observation
After acquiring the data cubes, we can calculate the bandpass images, given the throughput of each filter for a particular instrument. All calculations are performed in units of electron counts, enabling straightforward incorporation of PSF effects and instrumental and background noises. The electron counts for each pixel can be determined by providing the exposure time, , the number of exposures, , the effective area of the telescope, , and the pixel scale of the CCD camera, , as follows (see Section 9.2 in Ryon (2023)):
where and are Planck constant and the speed of light respectively; then, is the throughput of filter and is flux measured per wavelength, converted from flux measured per frequency in units of employed by SKIRT using the following equation:
The PSF describes how a point source of light, such as a star, is distributed across a detector, and is a fundamental characteristic of imaging system, as it determines the system’s ability to resolve the fine details. The PSF size is influenced by fundamental physical limits as well as practical effects such as atmospheric turbulence, optical aberrations, and detector effects. In ground-based observations, the most dominant factor is atmospheric turbulence, while in space-based ones, detector effects play a larger role. Furthermore, the PSF exhibits variations across the position of the CCD and observational pointing. Consequently, for simplicity, in our galaxy generation for specific instrument, we only considered the PSF kernels at an arbitrary position and convolved the ideal images with these kernels. We note that when convolving with the model PSF, the pixel sizes of the image and the kernel should be the same.
@@
where indicates the pixel of image on which the Poisson components are added; then, is the mean of expected electron counts in pixel, calculated as:
where is the pixels of image with PSF effects already applied; then, is the noise term that includes the contribution of two components, sky background noise, and dark current noise,
where and are sky background and dark current both in unit of electron count rate (). Between them, dark current, , comes from thermal activity within CCD, even in the absence of light, while the sky background can be calculated by (see Section 9.4 in Ryon (2023)):
where is the brightness of the sky background in Å-1. The sky brightness is contributed by several components, such as airglow, zodiacal light, light pollution, and terrestrial background (Ryon, 2023). These components can vary depending on factors like calendar day, observational pointing, orbit or site position, and atmospheric conditions. Therefore, we only considered brightness curves under arbitrary conditions for specific observational campaigns.
Subsequently, readout noise was introduced, which results from the random fluctuations during the process of reading out the charge from a CCD. This noise was modeled as Gaussian distribution with a mean of zero and a standard deviation corresponding to the readout level, , and is correlated with the number of exposures, ,
where represents the pixel of image with the readout noise added. Next, we subtracted the mean level of sky background and dark current from Equation 9 and expressed as
Finally, we obtained the galaxy images would be expected to be observed by specific instrument. The image unit in electron counts can be converted back to flux per frequency in . This procedure of including both the instrumental and sky background contributions ensures that the simulated data accurately represents the noisy conditions encountered in real-world astronomical observations. Alternatively, the overall noise level for the galaxy images can also be determined from the limiting magnitudes for a specific band in a given survey (Fortuni et al., 2023; Bottrell et al., 2019, 2024; Martin et al., 2022; Wilkinson et al., 2024; Euclid Collaboration et al., 2023). This approach is also implemented and detailed in Appendix C. Depending on the availability of certain survey parameters, either this approach or the one described above can be utilized.
@@ 3 Sample outputs
In this section, we showcase the generation capabilities of GalaxyGenius for a diverse range of photometric surveys conducted both in space and on ground. These surveys encompass notable ongoing and planned missions such as CSST, HST, JWST, Euclid, Roman, and HSC. Subsequently, we highlight the galaxies generated from EAGLE simulations.
@@ 3.1 CSST
The Chinese Space Station Telescope (CSST, Zhan (2018); Gong et al. (2019)) is a 2-meter aperture space telescope, planned to be launched at 2027. This telescope is equiped with cutting-edge instruments, including a spectrograph and a photometric system, enabling multi-wavelength observations and spectroscopic studies. The photometric system loads seven filters, , and covering the near-UV (NUV) to NIR wavelength range. The pixel scale for each filter is uniform of .
@@
Given that the filters of CSST barely reach NIR wavelengths, we utilized the ExtinctionOnly simulation framework to account for the extinction and scattering effects caused by dust components. Furthermore, we also employed the NoMedium mode for comparative purposes. The wavelength range in rest-frame was set to be from 0.1 to 1.2 micron to account for all filters. Then the wavelength grid was configured as linear with 300 bins. The exposure time, , is assumed to be 150 s, with number of exposure, , as four for the and bands and two for the other bands, according to the focal plane design. The sky backgrounds are calculated as , and based on the high sky background emission including earthshine and zodiacal light in optical wavelength coverage at earth orbit provided by HST 444https://hst-docs.stsci.edu/acsihb/chapter-9-exposure-time-calculations/9-7-tabular-sky-backgrounds. Other necessary instrumental parameters employed in mock observation are outlined in Appendix A. We set the viewing angle, inclination and azimuth, to be both 0, meaning that the line of sight is in parallel to the -axis. Additionally, for spiral galaxy, the viewing angles are set to be face-on and edge-on. These angles are calculated based on the total angular momentum for star particles within 30 kpc sphere (McAlpine et al., 2016).
@@
Three subhalos are presented as examples and their properties are tabulated in Table 2, including stellar mass, star formation rate (SFR), and the number of distinct particles, as discussed in Section 2.2. The ideal SEDs generated from SKIRT in redshifted wavelength range are illustrated in Figure 2. The optical depth at 5500Å is also calculated. For better comparison, deviations between different simulation modes are also displayed in the lower panel. The galaxy images in seven bands output by postprocess procedure introduced in Section 2.4 are depicted in Figure 3. The images are presented in logarithmic scale, and to mitigate the occurence of NaN values and enhance visual clarity, they are clipped at .
Subhalo 0 is an elliptical galaxy with a low SFR and a substantial number of BC03 particles. The slight deviation in the lower panel of SED for subhalo 0 (left panel of Figure 2) indicates that the low fraction of dust elements shown in Table 2 has a negligible impact on the SEDs in optical, irrespective of the simulation mode utilized, with the corresponding optical depth being . In contrast, subhalo 31 is an irregular galaxy characterized by a high SFR and a large number of particles belonging to MAPPINGS-III SED family, which are indicative of star-forming regions. The median panel of Figure 2 exhibits low continuum and several pronounced emission lines, such as H 6563, [SIII] and Pa 10941 from low to high wavelength. Furthermore, the dust exhibit substantial impacts on the continuum and emission lines, leading to substantial deviations in flux amplitude in optical wavelengths and yielding a high optical depth of 0.36. Subhalo 253881, on the other hand, is a spiral galaxy with a high SFR characterized by several of aforementioned emission lines. The continuum exhibits similar shape to subhalo 0, primarily due to the dominance of BC03 particles. As anticipated, the SED exhibits a significant reduction in the continuum for both face-on and edge-on directions, with attenuation more pronounced in the edge-on case. This is due to the increased dust column density encountered along the edge-on line of sight, resulting in an optical depth of 0.12 compared to 0.02 for the face-on view. These distinct behaviors demonstrate that the dust elements primarily attenuates emissions from star-forming regions, with the degree of attenuation increasing with the column density of dust through which the light propagates(Lu et al., 2022).
@@ 3.2 Euclid
The Euclid mission, developed by the European Space Agency (ESA), is a 1.2-meter space-based observatory designed to explore the nature of dark energy and dark matter by mapping the geometry of the Universe (Euclid Collaboration et al., 2024d). Lauched in 2023, Euclid operates at the second Lagrange point (L2) and is optimized for both visible and NIR observations.
@@
Euclid’s payload includes two main instruments: VISible Imaging Channel (VIS, Euclid Collaboration et al. (2024a)) and Near Infrared Spectrometer and Photometer (NISP, Euclid Collaboration et al. (2024c)). VIS covers the optical wavelength range from 500 to 900 nanometers with a pixel scale of , providing high-resolution, wide-field imaging. This instrument is designed for precise measurement of galaxy shapes, essential for weak gravitational lensing studies. While NISP operates in the NIR range of 920 to 2000 nm, including a photometric mode for imaging in three broad filters (Y, J, H) with pixel scales of , and a spectroscopic mode using a slitless grism. The spectroscopic mode is employed to measure redshifts of galaxies, rendering baryon acoustic oscillation (BAO) studies.
@@
We employed similar parameters to CSST (as mentioned in Section 3.1) to perform the simulation, and to cover the optical and NIR bands, the wavelength range are set to be 0.5 to 5 micron. Furthermore, DustEmission mode is activated to account for the radiative transfer process in IR bands. And for comparative purposes, ExtinctionOnly mode is also performed and analyzed. We employed dust recipe as Camps et al. (2016) (Equation 2) and dust model as Zubko et al. (2004) described in Section 2.2. The exposure time, , are configured as 600 s with one exposure. The PSFs are modeled using a Gaussian distribution with full width at half maximum (FHWM) as and for VIS, Y, J, and H bands respectively (Euclid Collaboration et al., 2024a, c). The sky backgrounds are calculated as 0.7162, 3.6863, 4.4556, and 3.8228 for these bands based on background models 555https://irsa.ipac.caltech.edu/applications/BackgroundModel/ considering that the telescope orbits at L2 point. Other necessary instrumental parameters are given in Appendix A.
The SEDs in the NIR are presented in Figure 2. We notice that the NIR SEDs closely follow those in the optical wavelengths, indicating that extinction and scattering are the dominant effects in both optical and NIR wavelength ranges. Additionally, the deviations between DustEmission and ExtinctionOnly are negligible, except for several emission lines at longer wavelength. This suggest that dust emission has a minimal impact within the IR wavelength range () considered in this study. Correspondingly, the galaxy images for the VIS and NISP bands are depicted in Figure 4. The NIR images reveal extended structures in comparison to the observations at optical bands.
@@ 3.3 HST WFC3
The Hubble Space Telescope (HST; Freedman et al. (2001); Momcheva et al. (2016)) is a 2.4-meter optical and UV space observatory launched in 1990, orbiting Earth at an altitude of approximately 540 km. HST operates across UV (115-320 nm), visible (320-760 nm), and NIR (760-2500 nm) wavelength ranges, providing high-resolution imaging and spectroscopy free from atmospheric distortion. The telescope’s pointing precision, fine guidance sensors, and advanced suite of instruments enable observations of faint and distance objects with unprecedented clarity.
The Wide Field Camera 3 (WFC3, Momcheva et al. (2016); Windhorst et al. (2011); O’Connor et al. (2024)), installed during the fifth servicing mission in 2009, is one of HST’s most versatile instruments. It features two independent optical channels: UVIS and IR channel. UVIS operates in 200-1000 nm wavelength range with a pixel CCD detector, optimized for UV and visible light imaging. This channel has a pixel scale of and a field of view of approximately ; meanwhile the IR channel covers 800-1700 nm with a HgCdTe array detector, with a pixel scale of , providing a field of view of . For more detailed information on instrumental design, we refer to the WFC3 official website 666https://www.stsci.edu/hst/instrumentation/wfc3/instrument-design/.
The PSF for each filter is taken from the standard PSFs described in (Anderson, 2016). The instrumental and sky background noises are calculated using the instrumental parameters and taking into account the operation of HST at Earth orbit. As an example, we considered several bands F275W, F390W, and F814W from UVIS channels and F105W, F110W, F125W, and F160W from the IR channels to illustrate the generation of galaxies from the TNG. The exposure time is similarly assumed to be 600 s with one exposure. The galaxy images in considered bands are depicted in Figure 5.